OpenStudy (anonymous):

Solve. 3x2 = 33x + 24. I need the exact quadratic formula answer and work thanks.

jimthompson5910 (jim_thompson5910):

\[\Large 3x^2 = 33x + 24\] \[\Large 3x^2 - 33x - 24 = 0\] Now use the quadratic formula to solve \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(-33)\pm\sqrt{(-33)^2-4(3)(-24)}}{2(3)}\] \[\Large x = \frac{33\pm\sqrt{1089-(-288)}}{6}\] \[\Large x = \frac{33\pm\sqrt{1377}}{6}\] \[\Large x = \frac{33+\sqrt{1377}}{6} \ \text{or} \ x = \frac{33-\sqrt{1377}}{6}\] \[\Large x = \frac{33+9\sqrt{17}}{6} \ \text{or} \ x = \frac{33-9\sqrt{17}}{6}\] \[\Large x = \frac{11+3\sqrt{17}}{2} \ \text{or} \ x = \frac{11-3\sqrt{17}}{2}\]

OpenStudy (anonymous):

thats what i got but these are the answers the question gives you...

jimthompson5910 (jim_thompson5910):

For some odd reason, they didn't fully simplify the radical term. I'm not sure why. In any event, the answer is choice A because sqrt(1377) = 3*sqrt(153) and this '3' will cancel.